# Spatial Knowledge Graph-Guided Multimodal Synthesis

### v1

@@ Abstract
Recent advances in multimodal large language models (MLLMs) have significantly enhanced their capabilities; however, their spatial perception abilities remain a notable limitation. To address this challenge, multimodal data synthesis offers a promising solution. Yet, ensuring that synthesized data adhere to spatial common sense is a non-trivial task. In this work, we introduce SKG2Data, a novel multimodal synthesis approach guided by spatial knowledge graphs, grounded in the concept of knowledge-to-data generation. SKG2Data automatically constructs a Spatial Knowledge Graph (SKG) to emulate human-like perception of spatial directions and distances, which is subsequently utilized to guide multimodal data synthesis. Extensive experiments demonstrate that data synthesized from diverse types of spatial knowledge, including direction and distance, not only enhance the spatial perception and reasoning abilities of MLLMs but also exhibit strong generalization capabilities. We hope that the idea of knowledge-based data synthesis can advance the development of spatial intelligence.
Spatial Knowledge Graph-Guided Multimodal Synthesis
Yida Xue1,  Zhen Bi1,2,  Jinnan Yang2,  Jungang Lou2, Huajun Chen1,  Ningyu Zhang1thanks:  Corresponding Author. 1Zhejiang University  2Huzhou University {xueyida, bizhen_zju, huajunsir, zhangningyu}@zju.edu.cn
1 Introduction
Despite significant advancements in Multimodal Large Language Models (MLLMs) Zhao et al. (2023b); Yin et al. (2023) for visual processing tasks, a critical limitation persists in their ability to comprehend spatial relationships Prasad et al. (2023); Liu et al. (2023b); Du et al. (2024); Lei et al. (2024); Tong et al. (2024); Shiri et al. (2024); Zheng et al. (2024). Tong et al. (2024) observe that MLLMs often operate with “eyes wide shut”, revealing fundamental deficiencies in visual-spatial comprehension even in state-of-the-art models. This limitation underscores a significant gap between MLLMs and human capabilities, as humans inherently excel in tasks requiring spatial intelligence.
@@ 1 Introduction
To address this challenge, multimodal data synthesis offers a promising solution, requiring the generation of large-scale, spatially coherent multimodal data that accurately adheres to real-world constraints. A straightforward approach involves developing computational models (or world models) capable of generating continuous synthetic data Du et al. (2024); Ray et al. (2024); Li et al. (2025). However, this method is inefficient due to redundant data across scenarios and demands frequent model updates when environments change, leading to significant maintenance costs. Notably, human spatial intelligence stems from the brain’s acquisition of spatial knowledge, which enables robust generalization through the application of cognitive frameworks Hodges et al. (1999); Herweg and Kahana (2018); Peer et al. (2021). This insight naturally leads to a new paradigm for multimodal data synthesis: modeling spatial knowledge to guide the targeted generation of informative data.
Inspired by this insight, as illustrated in Figure 1, we introduce SKG2Data, an innovative framework specifically designed to generate multimodal synthetic data for spatial understanding. The core idea is to leverage a specialized knowledge graph (KG) as a spatial knowledge representation to guide the generation of multimodal data. Specifically, we construct a Spatial KG (SKG) that simulates human-like understanding of spatial knowledge by characterizing scenes and modeling spatial relationships. SKG2Data utilizes the SKG, which encodes object attributes and their spatial relationships, to generate corresponding images and textual data.
@@
We automatically construct a multimodal instruction dataset for MLLMs based on the SKG, comprising 974 nodes and 4,731 triplets that encode spatial relationships. To ensure a fair comparison of results and prevent data contamination, we also use SKG2Data to synthesize an evaluation dataset, referred to as SKG2Data-Holdout, which includes 120 images and 566 single-choice questions. We fine-tune the LLaVA-1.6-Vicuna-7B Liu et al. (2024a) and Llama-3.2-Vision-11B Meta AI (2024) models using the synthesized data. The fine-tuned models show significant performance improvements on SKG2Data-Holdout and other spatial understanding benchmarks, while maintaining their general abilities. Additionally, we carry out comprehensive experiments to examine the effect of different spatial relation types, dataset sizes, and object densities on model performance.
@@
2 Preliminaries
Spatial Knowledge Graph.
In recent years, KGs have been widely applied across various domains, such as FACTKG Kim et al. (2023) for fact verification. Unlike traditional KGs, we introduce the Spatial KG (SKG), a specialized KG designed to encode spatial knowledge and guide multimodal data synthesis for spatial tasks. The SKG is defined as a KG where nodes represent objects and edges represent spatial relationships through triplets. Formally, SKG can be represented as:
@@ 2 Preliminaries
In contrast to Scene Graphs Zhu et al. (2017, 2022), which provide a general representation of scenes, the triplets in an SKG are specifically designed to emphasize spatial relations, with nodes $E$ containing detailed descriptions of objects. This structured representation enables the generation of complex and detailed multimodal synthesis data that emphasize spatial relationships.
Multimodal Data Synthesis.
The task of multimodal data synthesis involves generating high-quality input-output pairs related to images using a model $f$. For a given image $i$ (either pre-existing or synthesized), the task uses $i$, its description $d$, or a combination of both as inputs to $f$. The model $f$ then synthesizes a set of input $Q$ and their corresponding output $A$.
We focus on generating multimodal data, with an emphasis on capturing spatial relationships between objects. Our methodology involves the simultaneous creation of images and their corresponding question-answer pairs through a three-component framework: a multimodal generative model $f_{1}$, a text-to-image generative model $f_{2}$, and a SKG $G$. Inspired by LLM-grounded Diffusion (LMD) Lian et al. (2024), we use $f_{1}$ to generate bounding boxes $B$ and description $d$ under the guidance of $G$. These bounding boxes are then integrated into the layout-diffusion model $f_{2}$ to synthesize images. After image synthesis, $G$ and $B$ further assist $f_{1}$ in generating multiple question-answer pairs $(Q,A)$ relevant to the synthesized image $i$.
@@
For implementation, we use GPT-4o and GLIGEN Li et al. (2023) version of LMD as $f_{1}$ and $f_{2}$, respectively.
3 Spatial Knowledge Graph-Guided Data Synthesis
In this section, we introduce SKG2Data, a novel approach that takes advantage of a SKG that incorporates spatial positional relationships to guide the synthesis of multimodal data. As shown in Figure 2, SKG2Data consists of two core modules: Spatial KG Generation and Multimodal Data Synthesis.
3.1 Spatial Knowledge Graph Generation
Scenes and Objects Generation.
The real world encompasses a wide variety of scenes, each containing distinct sets of objects. To simulate the distribution of scenes and objects in real-world images, we leverage GPT-4o to generate a series of scenes and the objects likely to appear within them. Starting with a few scenes as few-shot examples, we use GPT-4o to generate an additional 160 safe, diverse, and realistic scenes, significantly enhancing the diversity of our dataset. For each scene, GPT-4o is further employed to produce a list of plausible objects that could exist in the given context. To improve object generation, particularly for uncommon scenes, we incorporate Wikipedia documents as external knowledge to augment GPT-4o’s understanding. This approach ensures that the synthesized objects exhibit distributions closely aligned with real-world images, effectively preventing the generation of incongruous or improbable objects within the specified scenes. Formally, we first employ $f_{1}$ to generate $m$ general scenes, denoted as $\{s_{1},s_{2},\dots,s_{m}\}$. For each scene $s_{i}$, we then generate $n$ objects that are likely to appear in that scene, expressed as:
@@ 3.1 Spatial Knowledge Graph Generation
Spatial Knowledge Graph Construction.
@@
3.2 Multimodal Data Synthesis
Image Data Generation.
End-to-end text-to-image generative models, such as Stable Diffusion Rombach et al. (2021) and Dall-E Betker et al. (2023), often encounter challenges in accurately controlling object positions and quantities. In contrast, layout-based diffusion models enable precise control over object placement and quantities, ensuring that the generated images closely conform to SKG. We employ GPT-4o to generate bounding boxes and captions under the guidance of SKG, and seamlessly incorporate these outputs into a GLIGEN-based model. Recognizing that diffusion models cannot guarantee perfection in every generated image, we generate multiple images for each SKG by varying the random seed.
@@ 3.2 Multimodal Data Synthesis
$\begin{gathered}B_{l}=\{b_{ij_{1}},b_{ij_{2}},\ldots,b_{ij_{l}}\}=f_{1}(E_{l},T_{l}),\\ \end{gathered}$ (7)
In order to synthesize data on whether an object exists in the image for subsequent tasks, we randomly select an object from other objects in current scene as a non-existent object $o_{neg}$, ensuring that this object is excluded from the synthesized image. The elements $E_{l}$, $T_{l}$, and $s_{i}$, along with the optional non-existent object $o_{neg}$, are input to $f_{1}$ to produce the caption $d$.
@@
Image Data Filtering.
During the image generation stage, a significant challenge is the potential for discrepancies, such as hallucinations, which can occur during the transformation of the SKG into bounding boxes and captions. Additionally, the image generation model may introduce its own inaccuracies. These issues can result in inconsistencies between the generated image and the original SKG. To address this challenge, we conduct an image validation process using GPT-4o as a image filtering agent. This process is aimed at evaluating and ensuring that the generated image aligns with the SKG. Formally, we utilize $f_{1}$ to assess the alignment of an image $i$ with the expected entities $E_{l}$ and textual descriptions $T_{l}$. If $f_{1}$ returns True, the image $i$ is preserved; otherwise, it is discarded. This process is represented by the following equation:
@@
After filtering, the dataset contains a total of 974 objects and 95 spatial relationships. The distribution of the top 15 objects and spatial relationships in the filtered images is illustrated in Figure 3.
@@
Table 1: Comparison with SoTA methods on five benchmarks.
Textual Data Generation.
@@
[…]
@@
Textual Data Filtering.
@@
Examples of the final training data generated are illustrated in Figure 7 (Appendix A).
4 Experiments
4.1 Experimental Settings
Baselines.
We conduct experiments on three open-source MLLMs: LLaVA-1.5-7B Liu et al. (2023a), LLaVA-1.6-7B Liu et al. (2024a), and Llama-3.2-Vision-11B Meta AI (2024). Both LLaVA-1.5 and LLaVA-1.6 use the Vicuna-7B Chiang et al. (2023) LLM as their LLM backbone.
Benchmarks.
For a comprehensive evaluation, we examine all models across two distinct task categories. The first category encompasses benchmarks that are either directly or partially related to spatial understanding. This includes the publicly available benchmark MMVP and COCO-Spatial Kamath et al. (2023). For COCO-Spatial, we utilize a subset named Two-obj that focuses on the positional relationships of two objects. We then convert the original positive and negative sample pairs into single-choice questions, which are more conducive to MLLMs testing. We leave a portion of the synthesized image to generate test questions called SKG2Data-Holdout, which are also used for evaluation. The second category involves general capability testing benchmarks for MLLMs, including MMStar and HallusionBench.
Implementation Details.
We fine-tune both LLaVA-v1.6 and Llama-3.2-Vision using training data. These models consist of a vision encoder, a projector, and an LLM backbone. During training, we tune the LLM backbone with LoRA modules  Hu et al. (2021), keeping the projector and vision encoder parameters fixed. The detailed settings are shown in Table 2 (Appendix A).
4.2 Main Results
Table 1 presents a comprehensive comparison of multiple models evaluated across five benchmarks, highlighting the strengths of SKG2Data.
Spatial Understanding Results.
Fine-tuned LLaVA-1.6 and fine-tuned Llama-3.2-vision demonstrate marked improvements across several challenging benchmarks for spatial tasks, highlighting the efficacy of our methods. On the KG2Data-Holdout benchmark, Fine-tuned LLaVA-1.6 achieves a score of 70.1(+1.5)↑, while Fine-tuned Llama-3.2-vision reaches 74.7(+1.4)↑, indicating robust performance enhancements. For the COCO-Spatial benchmark, Fine-tuned LLaVA-1.6 outperforms with a score of 79.3(+3.9)↑, and Fine-tuned Llama-3.2-vision impressively achieves 59.8(+13.9)↑, suggesting significant spatial understanding advancements. In the MMVP benchmark, Fine-tuned LLaVA-1.6 shows notable improvement with a score of 36.7(+4.7)↑, while Fine-tuned Llama-3.2-vision also makes gains, reaching 30.7(+1.4)↑. In summary, our experimental results indicate that synthesized training data can improve the spatial understanding ability of MLLMs.

### v2

@@ Abstract
Recent advances in Multimodal Large Language Models (MLLMs) have significantly enhanced their capabilities; however, their spatial perception abilities remain a notable limitation. To address this challenge, multimodal data synthesis offers a promising solution. Yet, ensuring that synthesized data adhere to spatial common sense is a non-trivial task. Our approach addresses this critical gap by providing a systematic framework for generating spatially coherent data. In this work, we introduce SKG2Data, a novel multimodal synthesis approach guided by spatial knowledge graphs, grounded in the concept of knowledge-to-data generation. SKG2Data employs an automated pipeline for constructing Spatial Knowledge Graph (SKG) that effectively captures human-like spatial cognition, including directional and distance relationships. These structured representations then serve as precise guidance for our integrated synthesis pipeline, where a diffusion model generates spatially-consistent images while a MLLM produces corresponding textual descriptions. The automated construction of SKG enables scalable generation of diverse yet realistic spatial configurations, overcoming the limitations of manual data collection and annotation. Extensive experiments demonstrate that data synthesized from diverse types of spatial knowledge, including direction and distance, enhance the spatial perception and reasoning abilities of MLLMs markedly, albeit with a slight cost to their general capabilities. We hope that the idea of knowledge-based data synthesis can advance the development of spatial intelligence [^1] Code is available at https://github.com/zjunlp/Knowledge2Data..
I Introduction
Despite significant advancements in Multimodal Large Language Models (MLLMs) [1, 2] for visual processing tasks, a critical limitation persists in their ability to comprehend spatial relationships[3, 4, 5, 6, 7, 8, 9]. Recent studies demonstrate that even state-of-the-art MLLMs operate with “eyes wide shut” [7], exhibiting fundamental deficiencies in visual-spatial comprehension. These models show marked inabilities to accurately estimate relative positions between objects, consistently fail in judging spatial distances, and possess a limited capacity for reasoning about the spatial relationships among multiple objects. This limitation underscores a significant gap between MLLMs and human capabilities, as humans inherently excel in tasks requiring spatial intelligence.
@@ I Introduction
To address this challenge, multimodal data synthesis offers a promising solution, requiring the generation of large-scale, spatially coherent multimodal data that accurately adheres to real-world constraints. A straightforward approach involves developing computational models (or world models) capable of generating continuous synthetic data [5, 10, 11]. However, this method proves inefficient due to redundant data across scenarios and the need for frequent model updates when environments change, resulting in substantial high computational costs for environment updates. In contrast, human spatial intelligence stems from the acquisition of spatial knowledge by the brain, which allows robust generalization through the application of cognitive frameworks [12, 13, 14]. This insight naturally leads to a new paradigm for multimodal data synthesis: modeling spatial knowledge to guide the targeted generation of informative data. Inspired by this insight, we introduce SKG2Data, an innovative framework specifically designed to generate multimodal synthetic data for spatial understanding as illustrated in Figure 1. The core idea is to leverage a specialized knowledge graph as a structured spatial representation to guide multimodal data generation. Specifically, we construct a Spatial KG (SKG) by employing GPT-4o to characterize objects and model their spatial relationships, thereby establishing a structured representation of spatial knowledge. The resulting SKG, which encodes object attributes and their spatial relationships, then serves as the key control mechanism in SKG2Data to direct the generation of corresponding images and textual data.
@@
We automatically synthesize SKG-based spatial multimodal data and design three ways to ensure the quality of synthetic data: human verification, external knowledge base enhancement, and automatic multimodal agent verification. After filtering, we construct a multimodal instruction dataset, which contains 974 nodes and 4,731 triplets encoding spatial relationships. To ensure a fair comparison of results and prevent data contamination, we also use SKG2Data to synthesize an evaluation dataset, referred to as SKG2Data-Holdout, which includes 120 images and 566 single-choice questions. We fine-tune the LLaVA[15] and Llama-3.2-Vision[16] models using the synthesized data. The fine-tuned models show performance improvements on SKG2Data-Holdout and other spatial understanding benchmarks, while maintaining their general abilities. Additionally, we carry out comprehensive experiments to examine the effect of different spatial relation types, dataset sizes, and object densities on model performance. Our work makes the following principal contributions:
  • We propose SKG2Data, a novel framework for generating multimodal synthetic data for spatial understanding tasks. It uses a SKG to guide the creation of high-quality multimodal data that reflects real-world spatial constraints, addressing the critical limitation of existing MLLMs in comprehending spatial relationships.
  • We automatically construct a multimodal instruction dataset and an evaluation dataset SKG2Data-Holdout based on the SKG2Data. These datasets are rich and diverse, providing comprehensive training data and a fair benchmark for assessing the spatial understanding capabilities of MLLMs.
  • We demonstrate the effectiveness of our synthesized data through extensive experiments. Fine-tuning MLLMs with our data significantly improves their performance on spatial understanding tasks while only slightly reducing their general capabilities. Our findings also reveal that directional knowledge is crucial for enhancing spatial perception and that increasing object density in the data improves fine-grained recognition.
II Related Work
II-A Multimodal Large Language Models
The remarkable success of Large Language Models (LLMs) [17, 18, 19] has driven significant advances in Multimodal Large Language Models (MLLMs) [20, 15, 21, 16, 22, 23, 24, 25]. These MLLMs typically employ transformer-based architectures with specialized visual encoders, such as ViT [26] or CLIP [27], that project visual inputs into the language model’s embedding space. This architectural innovation enables models to process and understand both text and visual information through a unified framework, achieving state-of-the-art performance on visual tasks.
The development of comprehensive evaluation benchmarks has been crucial for measuring progress in this field. General-purpose multimodal benchmarks like MMVet [28] and MMStar [29] assess broad capabilities including visual recognition, text generation, and commonsense reasoning. More specialized benchmarks like HallusionBench [30] and POPE [31] focus on identifying and quantifying hallucinations in MLLMs. Recent work has highlighted spatial reasoning as a particularly challenging capability for MLLMs. This has led to the creation of dedicated spatial understanding benchmarks [32, 5, 7] that systematically evaluate abilities like relative position estimation, distance judgment, and multi-object spatial relationships. These benchmarks reveal that while MLLMs excel at many visual-language tasks, their spatial comprehension still lags significantly behind human performance, particularly in complex real-world scenarios requiring spatial reasoning.
II-B Synthetic Data Generation
Synthetic data has garnered significant attention as a novel solution to the challenges of obtaining large, diverse, and high-quality datasets [33, 34, 35, 36], particularly in domains where real-world data collection is expensive, privacy-sensitive, or inherently limited. Multimodal data synthesis represents an advanced and rapidly evolving field with transformative potential across computer vision, natural language processing, and human-computer interaction applications. This paradigm involves the systematic creation of integrated multimodal data, combining text, images, audio, and sensor modalities through algorithmic generation processes. Specifically, for current MLLMs, many works focus on synthesizing data in both image and text modalities to train or test models.
Many existing studies primarily rely on pre-existing images [37, 38, 39], occasionally incorporating additional image information such as captions, bounding boxes, or OCR input [40, 41, 42]. However, another area of research transcends generating textual instruction data and focuses on creating high-quality images through various synthesis techniques rather than relying solely on existing image datasets. For example, REACHQA [43] and Multimodal Self-Struct [44] use code to precisely synthesize chart images. Additionally, Scenethesis [45] and EmbSpatial [5] utilize indoor 3D simulators to generate images, while SynCLR [46] and VisMin [47] employ diffusion models to produce new images. Furthermore, since LLMs tend to exhibit biases and follow a long-tailed distribution due to their lack of specialization in specific domains [48, 49], there is a growing body of research [50, 51, 52, 53, 54] that employs knowledge enhancement techniques, such as KG, to generate higher quality data.
In this work, we propose a method for multimodal data synthesis based on diffusion models and knowledge enhancement techniques. End-to-end text-to-image generative models, such as Stable Diffusion [55] and Dall-E [56], often encounter challenges in accurately controlling object positions and quantities. In contrast, layout-based diffusion models, such as GLIGEN-based models [57, 58], achieve accurate control over object placement and quantities by incorporating bounding boxes or explicit layout guidance. LLM-grounded Diffusion (LMD) [57] utilizes LLMs to generate bounding boxes as layout information for these layout-based diffusion models. Specifically, our method leverages spatial knowledge graphs to guide the generation of both images and corresponding textual descriptions, ensuring that the synthesized data adheres to real-world spatial constraints. Recent advances in diffusion models enable high-fidelity image generation, while knowledge graphs provide structured representations that maintain semantic and spatial coherence across modalities.
@@ II-B Synthetic Data Generation
III Preliminaries
III-A Spatial Knowledge Graph
In recent years, KGs have been widely applied across various domains, such as FACTKG [59] for fact verification. Unlike traditional KGs, we introduce the Spatial KG (SKG), a specialized KG designed to encode spatial knowledge and guide multimodal data synthesis for spatial tasks. The SKG is defined as a KG where nodes represent objects and edges represent spatial relationships through triplets. Formally, SKG can be represented as:
@@ III-A Spatial Knowledge Graph
In contrast to Scene Graphs [60, 61], which provide a general representation of scenes, the triplets in an SKG are specifically designed to emphasize spatial relations, with nodes $E$ containing detailed descriptions of objects. This structured representation facilitates the generation of complex, detailed multimodal data that accurately capture spatial relationships. By systematically modeling spatial relationships through SKG, our method bridges the gap between multimodal synthetic data and human-like spatial understanding.
III-B Multimodal Synthetic Data
The task of multimodal data synthesis involves generating high-quality input-output pairs related to images using a model $f$. For a given image $i$ (either pre-existing or synthesized), the task uses $i$, its description $d$, or a combination of both as inputs to $f$. The model $f$ then synthesizes a set of input $Q$ and their corresponding output $A$. We focus on generating multimodal data, with an emphasis on capturing spatial relationships between objects.
Our methodology involves the simultaneous creation of images and their corresponding question-answer pairs through a three-component framework: a multimodal generative model $f_{1}$, a text-to-image generative model $f_{2}$, and a SKG $G$. Inspired by LLM-grounded Diffusion (LMD) [57], we use $f_{1}$ to generate bounding boxes $B$ and description $d$ under the guidance of $G$. These bounding boxes are then integrated into the layout-diffusion model $f_{2}$ to synthesize images. After image synthesis, $G$ and $B$ further assist $f_{1}$ in generating multiple question-answer pairs $(Q,A)$ relevant to the synthesized image $i$. This process can be mathematically expressed as follows:
@@ III-B Multimodal Synthetic Data
For implementation, we use GPT-4o and GLIGEN [58] version of LMD as $f_{1}$ and $f_{2}$, respectively.
IV Spatial Knowledge Graph-Guided Data Synthesis
In this section, we introduce SKG2Data, a novel approach that takes advantage of a SKG that incorporates spatial positional relationships to guide the synthesis of multimodal data. As shown in Figure 2, SKG2Data consists of two modules: Spatial KG Generation and Multimodal Data Synthesis.
IV-A Spatial Knowledge Graph Generation
Scenes and Objects Generation
The real world encompasses a wide variety of scenes, each containing distinct sets of objects. To simulate the distribution of scenes and objects in real-world images, we leverage GPT-4o to generate a series of scenes and the objects likely to appear within them. Starting with a few scenes as few-shot examples, we use GPT-4o to generate an additional scenes. After human inspection and filtering, we finally obtain 160 safe, diverse, and realistic scenes, significantly enhancing the diversity of our dataset. For each scene, GPT-4o is further employed to produce a list of plausible objects that could exist in the given context. To improve object generation, particularly for uncommon scenes, we incorporate Wikipedia documents as external knowledge to ensure the trustworthiness and rationality of objects in each scene. This approach ensures that the synthesized objects exhibit distributions closely aligned with real-world images, effectively preventing the generation of incongruous or improbable objects within the specified scenes. Through observation of the generated objects, we find that most of the objects appear in high frequency in each real scene. Formally, we first employ $f_{1}$ to generate $m$ general scenes, denoted as $\{s_{1},s_{2},\dots,s_{m}\}$. For each scene $s_{i}$, we then generate $n$ objects that are likely to appear in that scene, expressed as:
@@ IV-A Spatial Knowledge Graph Generation
Spatial Knowledge Graph Construction
@@
Figure 4: Training data cases featuring diverse spatial relationships, synthesized through our proposed method.
@@
IV-B Multimodal Data Synthesis
Image Data Generation
We employ GPT-4o to generate bounding boxes and captions under the guidance of SKG, and seamlessly incorporate these outputs into a GLIGEN-based model. Recognizing that diffusion models cannot guarantee perfection in every generated image, we generate multiple images for each SKG by varying the random seed.
@@ IV-B Multimodal Data Synthesis
[…]
In order to synthesize data on whether an object exists in the image for subsequent tasks, we randomly select an object from other objects in current scene as a non-existent object $o_{neg}$, ensuring that this object is excluded from the synthesized image. The elements $E_{l}$, $T_{l}$, and $s_{i}$, along with the optional non-existent object $o_{neg}$, are input to $f_{1}$ to produce the caption $d$. The formula is as follows:
@@
Image Data Filtering
During the image generation stage, a significant challenge is the potential for discrepancies, such as hallucinations, which can occur during the transformation of the SKG into bounding boxes and captions. Additionally, the image generation model may introduce its own inaccuracies. These issues can result in inconsistencies between the generated image and the original SKG. To address this challenge, we conduct an image validation process using GPT-4o as a image filtering agent. This process is aimed at evaluating and ensuring that the generated image aligns with the SKG. Formally, we utilize $f_{1}$ to assess the alignment of an image $i$ with the expected entities $E_{l}$ and relationships $T_{l}$. If $f_{1}$ returns True, the image $i$ is preserved; otherwise, it is discarded. This process is represented by the following equation:
@@
By verifying the attributes and positional relationships of objects in the image through the agent, filter out images that do not match the information in SKG. After filtering, the dataset contains a total of 974 objects and 95 spatial relationships. The distribution of the top 15 objects and spatial relationships in the filtered images is illustrated in Figure 3.
TABLE I: Comparison of different models on five benchmarks.
@@
Examples of the training data generated are illustrated in Figure 4.
Textual Data Generation
@@
Entity-Based Data focuses on the existence, attributes, and quantities of objects in a given image. The construction of Entity-Based Data is primarily guided by the entity set $E_{l}$, along with explicitly defined non-existent objects $o_{neg}$. This type of data serves as general visual question-answering data, maintaining the general capabilities of MLLMs without requiring additional external datasets. In contrast, Relation-Based Data emphasizes the spatial relationships between objects. The generation of Relation-Based Data is based on the spatial relationship triplets $T_{l}$ and generated bounding boxes $B_{l}$. Relation-Based Data is specifically designed to improve the spatial reasoning capabilities of MLLMs. The generation process for each question $x$ and its corresponding answer $y$ is formalized as follows:
@@
Textual Data Filtering
@@
After filtering, data pairs that the GPT-4o cannot correctly answer solely by providing images can be filtered.
V Experiments
V-A Experimental Settings
Baselines
We conduct experiments on three open-source MLLMs: LLaVA-1.5-7B [40], LLaVA-1.6-7B [15], and Llama-3.2-Vision-11B [16]. Both LLaVA-1.5 and LLaVA-1.6 use the Vicuna-7B [62] LLM as their LLM backbone.
Benchmarks
For a comprehensive evaluation, we examine all models across two distinct task categories. The first category encompasses benchmarks that are either directly or partially related to spatial understanding. This includes the publicly available benchmark MMVP [7] and COCO-Spatial [32]. For COCO-Spatial, we utilize a subset named Two-obj that focuses on the positional relationships of two objects. We then convert the original positive and negative sample pairs into single-choice questions, which are more conducive to MLLMs testing. We leave a portion of the synthesized image to generate test questions called SKG2Data-Holdout, which are also used for evaluation. The second category involves general capability testing benchmarks for MLLMs, including MMStar [29] and HallusionBench [30].
Implementation Details
We fine-tune both LLaVA-v1.6 and Llama-3.2-Vision using training data. These models consist of a vision encoder, a projector, and an LLM backbone. During training, we tune the LLM backbone with LoRA modules [63], keeping the projector and vision encoder parameters fixed.
The experimental setup employs LoRA for fine-tuning with a global batch size of 64. Training occurs on NVIDIA A800 GPUs using a learning rate of $1\times 10^{-5}$. The LoRA configuration includes a rank ($r$) of 16, an alpha ($\alpha$) value of 32, and a dropout rate of 0.05. The models undergo training for 2 epochs with mixed precision enabled to optimize computational efficiency. These parameters ensure stable training while maintaining memory efficiency and model performance.
V-B Main Results
Table I presents a comprehensive comparison of multiple models evaluated across five benchmarks, highlighting the strengths of SKG2Data.
Spatial Understanding Results
Fine-tuned LLaVA-1.6 and fine-tuned Llama-3.2-vision demonstrate marked improvements across several challenging benchmarks for spatial tasks, highlighting the efficacy of our methods. On the KG2Data-Holdout benchmark, Fine-tuned LLaVA-1.6 achieves a score of 70.1(+1.5)↑, while Fine-tuned Llama-3.2-vision reaches 74.7(+1.4)↑, indicating robust performance enhancements. For the COCO-Spatial benchmark, Fine-tuned LLaVA-1.6 outperforms with a score of 79.3(+3.9)↑, and Fine-tuned Llama-3.2-vision impressively achieves 59.8(+13.9)↑, suggesting significant spatial understanding advancements. In the MMVP benchmark, Fine-tuned LLaVA-1.6 shows notable improvement with a score of 36.7(+4.7)↑, while Fine-tuned Llama-3.2-vision also makes gains, reaching 30.7(+1.4)↑. In summary, our experiments show that synthesized training data enhances MLLMs’ spatial understanding, even without scene or object generation tailored for these spatial datasets.